OpenStudy (anonymous):

The sun roof in Troy’s house is shaped like a rectangle. He increases the length of the sides of the sun roof to twice the size of the existing ones. How will this change affect the perimeter of the sun roof? It will be 4 times the original perimeter. It will be 16 times the original perimeter. It will be 2 times the original perimeter. It will be 8 times the original perimeter.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

If you double every dimension, you're doubling the perimeter...here's why P = 2(L + W) ... Start with the perimeter of a rectangle formula P = 2(2L + 2W) ... Double each dimension P = 2*2(L+W) ... Factor out 2 P = 4(L+W) ... Multiply Notice how we go from 2(L + W) to 4(L+W) So we doubled the original perimeter

OpenStudy (anonymous):

THANKS if i had unlimited medals they would all go to you :)

jimthompson5910 (jim_thompson5910):

lol thanks